# Inflation with vector fields revisited: heavy entropy perturbations and primordial black holes

### v2

@@ Abstract
We revisit inflation coupled with vector fields employing kinetic coupling in the comoving gauge. It is known that there is a cumulative effect $\tilde{I}N^{2}$ on curvature power spectrum. For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between vector fields and inflaton $\tilde{I}$ is not extremely small. In this paper, we explore the regime where $\tilde{I}\gg 1$, a realm that has not been extensively explored due to the limitations of perturbative methods. We found that the entropy perturbation becomes heavy in this regime and the cumulative effect decays away on super-horizon scales. Consequently, the power spectrum retains its scale invariance in the decoupling limit. By straightforwardly integrating out the heavy modes near horizon-crossing, we derive a low-energy effective field theory describing a massless adiabatic perturbation with an imaginary speed of sound $c_{s}^{2}=-1/3$. Namely, the inflation with vector fields presents a potential mechanism for generating primordial black holes.
@@ 1 Introduction
Inflationary physics has proven successful in phenomenologically explaining the large-scale structure of our universe. The simplest inflationary model, single-field inflation, predicts nearly scale-invariant, Gaussian, and adiabatic primordial perturbations. These predictions are confirmed by Cosmic Microwave Background (CMB) observations [1, 2, 3]. However, the single-field model encounters theoretical challenges. These inflationary models are expected to be effective field theories derived from UV completed theories, such as string theory. Despite the anticipated suppression of higher-dimension irrelevant operators by the UV cutoff, they play a crucial role at the low-energy scale during inflation. The UV sensitivity of inflation gives rise to issues like the $\eta$-problem [4, 5] and the super-Planckian problem of a single large field [6, 7, 5], disrupting the flatness of the scalar field potential.
@@
The study of introducing extra fields during inflation has attracted attention to evade the issues of single-field inflation. In particular, the study of multi-scalar fields exhibiting strongly non-geodesic motion has been extensively examined in recent years [8, 9, 10, 11, 12, 13, 14]. In contrast, our focus is on U(1) gauge fields, commonly known as vector fields, as additional fields during inflation. This choice stems not only from the ubiquity of gauge fields as essential components of the standard model (SM) and beyond but also from their rich phenomenological implications in cosmology [15, 16, 17, 18, 19]. The breaking of the conformal coupling between vector fields and gravity is necessary to generate sufficiently strong vector fields [19]. Analogous to the non-geodesic motion of scalar fields, vector fields should exhibit background dynamics and participate alongside scalar fields in driving inflation. The inflationary phase undergoes a transition to new attractors if the potential is steep enough [20], which is very similar to the hyperbolic inflation [8]. This kind of models was firstly constructed to explain the statistical anisotropy in the CMB [21, 22, 23, 24] (or read the review [25]). The vector field is coupled to gravity through $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$. After suitable choices of $f(\phi)$, the dynamics of background vector fields provides an effective potential for the slow-roll scalar field.
@@
Under this new attractor, the power spectrum of curvature perturbation is sourced by the vector field, yielding an additional contribution
$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)IN_{k}^{2}\right],\ \ \ \ \ \ (I\ll 1)$ (1.1)
@@
where $N_{k}\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed mode elapsed since horizon-crossing, $\tau_{e}$ is time of end of the inflation. $I$ is the parameter related to specific models. For one U(1) vector field, $\Theta(\theta)=24\sin^{2}\theta$, where $\theta$ is the angle between mode and direction of the vector field [22, 26]. The term $IN_{k}^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $I<10^{-7}$ for a sufficiently lengthy inflationary period $N_{k}\sim 60$ [22, 26].The exceedingly small parameter implies the presence of a very small but non-zero vector field, potentially giving rise to fine-tuning challenges [27, 28]. Studies have revealed that employing multiple vector fields can mitigate the anisotropy, eventually converging to an isotropic attractor [24]. For the isotropic configuration of a triad of vector fields, $\Theta(\theta)=16$, relaxing the constraint on $I$ [29]. However, it still needs to be constrained to $I<10^{-5}$ due to the tilt of the power spectrum caused by the e-folding number $N$. This is a cumulative effect, i.e., the curvature is sourced by non-zero vector fields and evolves on super-horizon scales.
@@
However, this does not imply that results for $I\sim\mathcal{O}(1)$ are undesirable. If we require vector fields to handle a sufficiently large field displacement, indicating that $I$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates the adiabatic and scale-invariant initial conditions. However, this result lacks reliability because the perturbative method breaks down in this regime. It has been demonstrated that for $I\gtrsim 0.2$, the calculation by using in-in formalism starts to deviate from the numerical results and it is difficult to track the dynamics near and outside the horizon analytically [30, 31]. Numerical findings indicate that the power spectrum does not actually evolve on super-horizon scales [31, 20]. In this paper, we revisit this dilaton-gauge coupling model in comoving gauge and thoroughly explore the significant $I$ case with the assistance of effective field theory (EFT). We will show, the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation. $IN_{k}^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on one parameter. This enhancement also provides a mechanism for generating primordial black holes.
@@ 2 Inflation with vector fields
We are considering the spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$, where ˙$\dot{}\equiv d/dt$ is cosmological time derivative. The Arnowitt-Deser-Misner(ADM) formalism of the metric is [32]
@@
where $N$ is the lapse function and we choose the cosmological time $N=1$ for background, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of hypersurface. We consider the model of inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
@@
where $g$ is the determinant of metric, $V(\phi)$ is the inflationary potential and $f_{ab}$ is the kinetic coupling and only depends on the inflaton. The strength of vector is $F^{a}_{\ \mu\nu}=\partial_{\mu}A^{a}_{\ \nu}-\partial_{\nu}A^{a}_{\ \mu}$.
@@ 2.1 Background
We consider one dilaton field $\phi$ as inflaton and a triplet of $U(1)$ fields and take an isotropic configuration of background evolution, which can be realized by the choice of $U(1)$ fields as following [33, 34]
@@
The identification of the spatial indices and internal indices of the field space allows us to achieve the internal gauge transformation of three $U(1)$ gauge fields by the $O(3)$ rotation of the three spatial dimensions in real space. We mention that the discussion doesn’t rely on the number of the $U(1)$ fields. If we have only one $U(1)$ field and kinetic coupling with dilaton, there may be anisotropy left in our universe when the energy density of $U(1)$ gauge field is not so small [22, 26]. Although we will see, the theoretical parameterization of power spectrum is different when this energy density is large enough, we still consider isotropic case so that we can only treat scalar modes at linear order.
@@
where $\epsilon\equiv-\dot{H}/H^{2}$, $\epsilon_{\phi}\equiv\dot{\phi}^{2}/(2M_{\text{pl}}^{2}H^{2})$ and $\epsilon_{A}\equiv f^{2}\dot{\mathbb{A}}^{2}/(a^{2}M_{\text{pl}}^{2}H^{2})$. It’s also useful to define $\dot{\sigma}^{2}\equiv\dot{\phi}^{2}/2+f^{2}\dot{\mathbb{A}}^{2}/a^{2}$ as the “kinetic” energy of the matter fields. Then another gravitational equation can be written as
@@
where $V_{\phi}\equiv\partial V/\partial\phi$ and $f_{\phi}\equiv\partial f/\partial\phi$. In this paper we assume that $\dot{\phi}<0$ and $\dot{\mathbb{A}}>0$. We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and dilaton [^1] The parameter $I$ in [29] or in (1.1) is related to $h$ in our paper as $I\simeq 2h^{2}$ for $h\ll 1$.
@@
But here the ratio $h$ needn’t to be a small parameter. In fact, we will consider the larger $h\gg 1$ on small scales for PBH’s production. We will also define the slow-roll parameter $\epsilon_{h}\equiv\dot{h}/(Hh)$, $\eta_{h}\equiv\dot{\epsilon}_{h}/(H\epsilon_{h})$, $\eta_{\phi}\equiv\ddot{\phi}/(H\dot{\phi})$ and $\eta_{A}\equiv\dot{\epsilon}_{A}/(H\epsilon_{A})$. Taking time derivative on the Friedmann equation one can obtain
@@ 2.2 Quadratic action in comoving gauge
Thanks to the isotropic background and we will work in the quadratic action and the linear fluctuations of the theory hence the vector and tensor fluctuations are decoupled from scalar ones at linear order. After using the helicity decomposition the scalar parts of the $U(1)$ fields can be written as
@@
We do not distinguish indices $a$ and $i$ here because we identified the spatial rotation symmetry with the internal global $O(3)$ symmetry of the space of the gauge fields. The gauge fields enjoys local $U(1)$ symmetry $A^{a}_{\ \mu}\rightarrow A^{a}_{\ \mu}+\partial_{\mu}\rho^{a}$, where $\rho^{a}$ is a arbitrary function. One can fix the scalar gauge freedom of $U(1)$ fields by $\mathbb{M}=0$ [20]. The $\mathbb{Y}$ is the constrain d.o.f and we will eliminate it latter after fixing the gauge freedoms of gravity.
The scalar fluctuations of gravity and scalar fields can be written as
@@
There are four gauge freedoms of the time and spatial coordinate choice. If we have scalar field, in addition to vector fields we still have one physical d.o.f in the theory. The spatial flat gauge is usually used to study the dilaton-gauge theory. In this gauge, this physical d.o.f is the fluctuation of the scalar field $\delta\phi$. On the other hand, the “comoving” gauge is also used in the two-scalar fields [35, 36]. In this gauge, the adiabatic fluctuation is set to zero. The physical d.o.f are the curvature and entropy fluctuations.
In this paper we adopt the similar comoving gauge for the dilaton-gauge model. The adiabatic fluctuation is given by $Q_{\sigma}\equiv-\sqrt{2}\dot{\sigma}\delta u$, where $\delta u$ is the velocity potential given by $\delta T^{0}_{i}\equiv(\rho+p)\partial_{i}\delta u$. For the isotropic one-form gauge fields, we have $\delta T^{0}_{\ i}=-\dot{\phi}\partial_{i}\delta\phi-(2f^{2}\dot{\mathbb{A}}/a^{2})\partial_{i}\delta\mathbb{A}$ [20]. Here $\rho$ and $p$ are the total energy density and pressure of the system and in our model we have $\rho+p=\dot{\phi}^{2}+2f^{2}\dot{\mathbb{A}}^{2}/a^{2}=2\dot{\sigma}^{2}$. It’s useful to define the variable $\delta Q=\sqrt{2}f\delta\mathbb{A}/a$. Then the adiabatic fluctuation is given by
@@
This is the unitary gauge in the gauge theory, whose Nambu-Goldstone boson is eaten by the metric, as we will see in the next section. Then we have fixed all gauge freedom of the theory. The d.o.f $\mathbb{U}$ in the vector fields is the magnetic fluctuation and decouple with any other modes in our configuration. It only contribute to the isocurvature fluctuation hence we ignore it in the following discussion. We also define the entropy fluctuations. Defining [37, 29]
@@
which is similar to the entropy fluctuation in the two-scalar fields. But here the entropy fluctuations is not equivalent to the isocurvature one because we still have another contribution $\mathbb{U}$, which is decoupled with $\zeta$ and $\mathcal{F}$.
After imposing the comoving gauge, the equation of motion of the non-dynamical d.o.f $\alpha$ and $\mathbb{Y}$ can be derived and the solutions read
$\alpha=$ $\frac{\dot{\zeta}}{H},$ (2.16)
$\partial^{2}\mathbb{Y}=$ $\dot{\mathbb{A}}\left(\zeta-\frac{\dot{\zeta}}{H}\right)-\frac{a}{\sqrt{2}f}\left(\frac{\mathcal{F}}{\sqrt{1+2h^{2}}}\right)^{.}$
$-\frac{a}{\sqrt{2}f}\left(H-\frac{\dot{f}}{f}+4f_{\phi}\frac{\dot{\mathbb{A}}}{a}h\right)\frac{\mathcal{F}}{\sqrt{1+2h^{2}}}.$ (2.17)
The d.o.f $\beta$ is totally eliminated from the quadratic action when we use the constrain equation of $\alpha$ hence we don’t need to write down its specific expression. Then after using the background equations (2.6) and (2.9) to eliminate the $V$, $V_{\phi}$ and $f_{\phi}$, and inserting the constrain equations into the quadratic Lagrangian, we obtain
@@
where we have discarded slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$. This is the decoupling limit $\epsilon\to 0$ in the EFT. Then $\epsilon_{\phi}$ and $\epsilon_{A}$ are also vanishing because they are positive quantities.
We find that the curvature and entropy fluctuations are both massive. The mass term of $\zeta$ is only dependent on $h$. For the mass term of $\mathcal{F}$, the second term can be regards as the contribution from the curved field space, whose curvature is given by $R_{\text{fs}}=-f_{\phi\phi}/f$. If we assume $h=0$ initially, where we are in the single-field inflation regime, the entropy mass $m_{s}^{2}$ is still non-vanishing and have large negative contribution if the curvature of field space is negatively large enough. In the slow-roll regime, the second derivative of $f$ in the entropy mass becomes $f_{\phi\phi}/f\simeq f_{\phi}^{2}/f^{2}$. Then using (2.9) the mass square of the entropy fluctuation reads
@@
Therefore the entropy fluctuation has imaginary mass. This is similar to the so-called geometrical destabilization [11, 12], which is firstly studied in the multi-scalar inflation and then in the inflation with multiple vector fields [20]. The differences are that the curvature fluctuation $\zeta$ is massive and we have an additional coupling $\zeta\mathcal{F}$. Hence we next investigate the mode solutions of this Lagrangian.

### v3

@@ Abstract
We revisit inflation coupled with vector fields employing kinetic coupling in the comoving gauge. It is known that there is a cumulative effect $IN^{2}$ on the curvature power spectrum. For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between the vector fields and the inflaton $I$ is not extremely small. In this paper, we explore a regime where $I\gg 1$. This regime has not been extensively explored due to the limitations of perturbative methods. We found that the entropy perturbation becomes heavy in this regime and the cumulative effect decays away on super-horizon scales. Consequently, the power spectrum retains its scale invariance in the decoupling limit. By straightforwardly integrating out the heavy modes near horizon-crossing, we derive a low-energy effective field theory describing a massless adiabatic perturbation with an imaginary speed of sound $c_{s}^{2}=-1/3$. Namely, the inflation with vector fields presents a potential mechanism for generating primordial black holes.
@@ 1 Introduction
Inflationary physics has proven to be successful in phenomenologically explaining the large-scale structure of our universe. The simplest inflationary model, single-field inflation, predicts nearly scale-invariant, Gaussian, and adiabatic primordial perturbations. These predictions are confirmed by Cosmic Microwave Background (CMB) observations [1, 2, 3]. However, the single-field model encounters theoretical challenges. These inflationary models are expected to be effective field theories derived from UV completed theories, such as string theory. Despite an anticipated suppression of higher-dimension irrelevant operators by UV cutoff, they play a crucial role at the low-energy scale during inflation. The UV sensitivity of inflation gives rise to issues like the $\eta$-problem [4, 5] and the super-Planckian problem of a single large field [6, 7, 5], disrupting the flatness of the scalar field potential.
@@
Study of introducing extra fields during inflation has attracted attention to evade these issues of single-field inflation. In particular, the study of multi-scalar fields exhibiting strongly non-geodesic motion has been extensively examined in recent years [8, 9, 10, 11, 12, 13, 14]. In contrast, our focus is on U(1) gauge fields, commonly known as vector fields, as additional fields during inflation. This choice stems not only from ubiquity of gauge fields as essential components of the standard model (SM) and beyond but also from their rich phenomenological implications in cosmology [15, 16, 17, 18, 19]. Breaking of conformal couplings between vector fields and gravity is necessary to generate sufficiently strong vector fields [19]. Analogous to the non-geodesic motion of scalar fields, vector fields should exhibit background dynamics and participate alongside the scalar fields in driving inflation. The inflationary phase undergoes a transition to new attractors if the potential is steep enough [20], which is very similar to the hyperbolic inflation [8]. This kind of models was firstly constructed to explain the statistical anisotropy in the CMB [21, 22, 23, 24] (or read the review [25]). The vector field is coupled to gravity through $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$. After suitable choices of $f(\phi)$, dynamics of background vector fields provides an effective potential for the slow-roll scalar field.
@@
Under this new attractor, the power spectrum of the curvature perturbation is sourced by the vector fields, yielding an additional contribution
$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)\mathcal{I}N_{k}^{2}\right],\ \ \ \ \ \ (\mathcal{I}\ll 1)$ (1.1)
@@
where $N_{k}\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed modes elapsed since horizon-crossing, $\tau_{e}$ is the end time of the inflation. $\mathcal{I}$ is a parameter that related to specific models. For one U(1) vector field, $\Theta(\theta)=24\sin^{2}\theta$, where $\theta$ is the angle between mode and direction of the vector field [22, 26]. The term $\mathcal{I}N_{k}^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $\mathcal{I}<10^{-7}$ for a sufficiently lengthy inflationary period $N_{k}\sim 60$ [22, 26]. The exceedingly small parameter implies the presence of a tiny but non-zero vector field, potentially giving rise to fine-tuning challenges [27, 28]. Studies have revealed that employing multiple vector fields can mitigate the anisotropy, eventually converging to an isotropic attractor [24]. For the isotropic configuration of a triad of vector fields, $\Theta(\theta)=16$, relaxing the constraint on $\mathcal{I}$ [29]. However, it still needs to be constrained to $\mathcal{I}<10^{-5}$ due to the tilt of the power spectrum caused by the e-folding number $N_{k}$. This is a cumulative effect, i.e., the curvature is sourced by the non-zero vector fields and evolves on super-horizon scales.
@@
However, this does not imply that results for $\mathcal{I}\sim\mathcal{O}(1)$ are undesirable. If we require vector fields to handle a sufficiently large field displacement, indicating that $\mathcal{I}$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates adiabatic and scale-invariant initial conditions. However, this result lacks reliability because the perturbative method breaks down in this regime. It has been demonstrated that for $\mathcal{I}\gtrsim 0.2$, the calculation by using the in-in formalism starts to deviate from the numerical results and it is difficult to track the dynamics near and outside the horizon analytically [30, 31]. Numerical findings indicate that the power spectrum does not actually evolve on super-horizon scales [31, 20]. In this paper, we revisit this dilaton-gauge coupling model in the comoving gauge and thoroughly explore the significant $\mathcal{I}$ case with the assistance of the effective field theory (EFT). We will show, that the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation. $\mathcal{I}N_{k}^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on one parameter. This enhancement also provides a mechanism for generating primordial black holes.
@@ 2 Inflation with vector fields
We are considering a spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$, where ˙$\dot{}\equiv d/dt$ isthe cosmological time derivative. The Arnowitt-Deser-Misner(ADM) formalism for the metric is [32]
@@
where $N$ is the lapse function and we choose the cosmological time $N=1$ for background, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of the hypersurface. We consider a model of an inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
@@
where $g$ is the determinant of the metric, $V(\phi)$ is the inflationary potential and $f_{ab}$ is the kinetic coupling, which only depends on the inflaton. The strength of the vector fields is $F^{a}_{\ \mu\nu}=\partial_{\mu}A^{a}_{\ \nu}-\partial_{\nu}A^{a}_{\ \mu}$.
@@ 2.1 Background
The configuration of the gauge fields is chosen for preserving isotropy [33, 34]. One can consider a large number of randomly vector fields. These vector fields have about the same magnitude of order $A$ initially. Then, for example, [34]
$\sum_{a}^{N\gg 1}A^{a}_{\ i}A^{a}_{\ j}\simeq\frac{N}{3}A^{2}\delta_{ij}+\mathcal{O}(1)\sqrt{N}A^{2}$ (2.3)
is dominated by the $\delta_{ij}$ parts. The corrections proportional to $\sqrt{N}$ are due to stochastic random distribution of directions of the fields and they do not vanish for $i\neq j$. Moreover, if the coupling $f_{ab}$ are uniform, the isotropic configuration is an attractor in the phase space for the number of vector fields greater than two. This has been confirmed numerically in [24]. If the field is a SU(2) non-Abelian gauge field, it’s natural to have this uniform coupling. The gauge coupling can be ignored during inflation thus SU(2) gauge field looks exactly like a triplet of U(1) gauge fields [35].
Therefore in this paper we assume one dilaton field $\phi$ as the inflaton and a triplet of $U(1)$ fields and take an isotropic configuration of the background evolution, which can be realized by the following choice of $U(1)$ fields [36, 37]
@@
The identification of the spatial indices and the internal indices of the field space allows us to achieve an internal gauge transformation of the three $U(1)$ gauge fields by a $O(3)$ rotation of the three spatial dimensions in real space. We note that the discussion doesn’t rely on the number of the $U(1)$ fields. If we have only one $U(1)$ field and kinetic coupling with dilaton, there may be anisotropy left in our universe when the energy density of the $U(1)$ gauge field is not so small [22, 26].
@@
where $\epsilon\equiv-\dot{H}/H^{2}$, $\epsilon_{\phi}\equiv\dot{\phi}^{2}/(2M_{\text{pl}}^{2}H^{2})$ and $\epsilon_{A}\equiv f^{2}\dot{\mathbb{A}}^{2}/(a^{2}M_{\text{pl}}^{2}H^{2})$. It’s also useful to define $\dot{\sigma}^{2}\equiv\dot{\phi}^{2}/2+f^{2}\dot{\mathbb{A}}^{2}/a^{2}$ as the “kinetic” energy of the matter fields. Another gravitational equation can then be written as
@@
where $V_{\phi}\equiv\partial V/\partial\phi$ and $f_{\phi}\equiv\partial f/\partial\phi$. In this paper we assume that $\dot{\phi}<0$ and $\dot{\mathbb{A}}>0$. We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and the dilaton [^1] The parameter $\mathcal{I}$ in [29] or in (1.1) is related to $h$ in our paper as $\mathcal{I}\simeq 2h^{2}$ for $h\ll 1$.
@@
But here, the ratio $h$ needn’t to be a small parameter. In fact, we will consider the larger $h\gg 1$ on small scales for PBH’s production. We will also define the slow-roll parameter $\epsilon_{h}\equiv\dot{h}/(Hh)$, $\eta_{h}\equiv\dot{\epsilon}_{h}/(H\epsilon_{h})$, $\eta_{\phi}\equiv\ddot{\phi}/(H\dot{\phi})$ and $\eta_{A}\equiv\dot{\epsilon}_{A}/(H\epsilon_{A})$. Taking the time derivative of the Friedmann equation one obtains
@@ 2.2 Quadratic action in comoving gauge
Thanks to the isotropic background and we will work in the quadratic action and the linear fluctuations of the theory. Hence, the vector and tensor fluctuations are decoupled from the scalar ones at linear order. After using the helicity decomposition the scalar parts of the $U(1)$ fields can be written as
@@
We do not distinguish between the indices $a$ and $i$ here because we identified the spatial rotation symmetry with the internal global $O(3)$ symmetry of the space of the gauge fields.
The scalar fluctuations of gravity and the scalar field can be written as
@@
We have analysed the vector and tensor perturbations in our previous work [20]. The $U$ and the vector perturbations are massless (see (5.10) of [20]). Therefore, there is no instability in these perturbations so we didn’t discuss therm in this work. The tensor perturbations are almost the same as the single-filed inflation thus they are not interesting for our discussion (see (5.14) of [20]).
We have a total of nine scalar degrees of freedom, where four from the metric and five from the matter sector. There are four gauge freedoms of the time and spatial coordinate choice. If we have one scalar field, there are two scalar gauge freedoms in the theory. On the other hand, if a triple of gauge fields is considered, it enjoys a local $U(1)$ symmetry $A^{a}_{\ \mu}\rightarrow A^{a}_{\ \mu}+\partial_{\mu}\rho^{a}$, where $\rho^{a}$ is an arbitrary function. This allows for another one gauge freedom. Fixing each gauge freedom removes one constraint. In conclusion, we can eliminate $2\times(2+1)=6$ d.o.f, leaving $3$ remaining.
One can fix the scalar gauge freedom of the $U(1)$ fields by $\mathbb{M}=0$ [20]. The spatial flat gauge is usually used to study the dilaton-gauge theory. In this gauge, the physical d.o.f are the fluctuation of the scalar field $\delta\phi$. On the other hand, the “comoving” gauge is also used in the two-scalar fields [38, 39]. In this gauge, the adiabatic fluctuation is set to zero. The physical d.o.f are the curvature and entropy fluctuations.
In this paper we adopt a similar comoving gauge for the dilaton-gauge model. The adiabatic fluctuation is given by $Q_{\sigma}\equiv-\sqrt{2}\dot{\sigma}\delta u$, where $\delta u$ is the velocity potential given by $\delta T^{0}_{i}\equiv(\rho+p)\partial_{i}\delta u$. For the isotropic one-form gauge fields, we have $\delta T^{0}_{\ i}=-\dot{\phi}\partial_{i}\delta\phi-(2f^{2}\dot{\mathbb{A}}/a^{2})\partial_{i}\delta\mathbb{A}$ [20]. Here $\rho$ and $p$ are the total energy density and pressure of the system, and in our model, we have $\rho+p=\dot{\phi}^{2}+2f^{2}\dot{\mathbb{A}}^{2}/a^{2}=2\dot{\sigma}^{2}$. It’s useful to define a variable $\delta Q=\sqrt{2}f\delta\mathbb{A}/a$. Then the adiabatic fluctuation is given by
@@
This is the unitary gauge in the gauge theory, whose Nambu-Goldstone boson is eaten by the metric. Then we have fixed all gauge freedom of the theory. The d.o.f $\mathbb{U}$ in the vector fields is the magnetic fluctuation and decouples from any other modes in our configuration. It only contributes to the isocurvature fluctuation hence we ignore it in the following discussion. We also define the entropy fluctuations. Defining [40, 29]
@@
which is similar to the entropy fluctuation in systems of two-scalar fields. But here the entropy fluctuations is not equivalent to the isocurvature one because we still have another contribution $\mathbb{U}$, which is decoupled from the $\zeta$ and $\mathcal{F}$.
After imposing the comoving gauge, the solutions of the equations of motion of the non-dynamical d.o.f $\alpha$, $\beta$ and $\mathbb{Y}$ can be derived (see Appendix A). Then after using the background equations (2.7) and (2.10) to eliminate the $V$, $V_{\phi}$ and $f_{\phi}$, and inserting the constrain equations into the quadratic Lagrangian, we obtain
@@
where we have discarded slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$. This is the decoupling limit $\epsilon\to 0$ so that the higher-order terms of $\epsilon$ are discarded. The higher order of $\epsilon_{\phi}$ and $\epsilon_{A}$ are also vanishing because they are positive quantities as we have $\epsilon=\epsilon_{\phi}+\epsilon_{A}$.
We find that the curvature and entropy fluctuations are both massive. The mass term of $\zeta$ is only dependent on $h$. For the mass term of $\mathcal{F}$, the second term can be regarded as the contribution from the curved field space, whose curvature is given by $R_{\text{fs}}=-f_{\phi\phi}/f$. If we assume $h=0$ initially, where we are in the single-field inflation regime, the entropy mass $m_{s}^{2}$ is still non-vanishing and have large negative contribution if the curvature of field space is negatively large enough. In the slow-roll regime, the second derivative of $f$ in the entropy mass becomes $f_{\phi\phi}/f\simeq f_{\phi}^{2}/f^{2}$. Then using (2.10) the mass square of the entropy fluctuation reads
@@
Therefore the entropy fluctuation has an imaginary mass. This is similar to the so-called geometrical destabilization [11, 12], which is firstly studied in the multi-scalar inflation and then in the inflation with multiple vector fields [20]. The differences are that the curvature fluctuation $\zeta$ is massive, and we have an additional coupling $\zeta\mathcal{F}$. Hence we next investigate the mode solutions of this Lagrangian.